Describe the sampling distribution of p. Assume the size of the population is 20,000.

n=700​, p=0.5

Determine the mean of the sampling distribution of

mu Subscript ModifyingAbove p with caret Baseline equals

μp=____

​(Round to one decimal place as​ needed.)

Determine the standard deviation of the sampling distribution of

sigma Subscript ModifyingAbove p with caret Baseline equals

σp=____

​(Round to three decimal places as​ needed.)
Answer #1

Solution

Given that,

p = 0.5

1 - p = 1 - 0.5 = 0.5

n = 700

The mean of the sampling distribution of proportion is ,

= p = 0.5

The standard deviation of the sampling distribution is ,  

=  p ( 1 - p ) / n

=   (0.5 * 0.5) / 700 = 0.019

= 0.019